I really need your help here is my question
You have made the scores of 81,76 and 82 on three quizzes. What is the least you can make on the fourth quiz to have an average of at least 80?

Answer by bam878s(77)

Average=(quiz1+quiz2+quiz3+quiz4)\4
So to get an 80 we solve
80=(81+76+82+x)\4 for x (let x=quiz4)
multiply both sides by 4
320=81+76+82+x
320=239+x subtract 239 from both sides
320-239=239+x-239
81=x
You need to make at least an 81.